At 298 K, a 0.100 M solution of acetic acid is 1.34% ionised. Calculate the ionisation
constant of acetic acid.
1
Expert's answer
2016-08-30T04:21:04-0400
Ka = ɑ^2*C/(1 - ɑ); ɑ = 1.34% or 0.0134; C = 0.1 M.
Ka = 0.01342 × 0.1 / (1 - 0.0134) = 1.8 × 10^{-5}.
Answer: 1.8 × 10^{-5}.